OpenStudy (gracygirl):

Help please! Maple syrup worth $6.00 a gallon and corn syrup worth .80 cents a gallon are used to make a mixture worth $2.36 a gallon. How many gallons of each kind of syrup are need to make 50 gallons of the mixture?

OpenStudy (gracygirl):

I'm not how to solve this one :/

OpenStudy (gracygirl):

@AloneS @Awolflover1 @Frozen15

OpenStudy (frozen15):

here is what I did.. Maple syrup worth $6.00 a gallon and corn syrup worth .80 cents a gallon are used to make a mixture worth $2.36 a gallon. How many gallons of each kind of syrup are need to make 50 gallons of the mixture? {maple syrup: gallons, corn syrup: gallons} let quantity of maple syrup used be x gallons corn syrup used will be 50-x gallons 6*x+0,8(50-x)= 2.36*50 6x+40-0.8x=118.00 5.2x=118-40 5,2x=74 x=74/5.2 x=14.2 gallons 35.8 gallons will be corn syrup. So therefore , the amount of Corn Syrup is your answer you are looking for

OpenStudy (gracygirl):

Okay thank you! Each time it gets a little less confusing :P

OpenStudy (frozen15):

your welcome haha , I did get this solution by "algebra.com" but I understood it and gave it to you for information of each step and such

OpenStudy (gracygirl):

Okay :)

OpenStudy (frozen15):

if you need anymore help , tag me in a new post or pm me :)